# Trade-offs in Data Memorization via Strong Data Processing Inequalities

### v1

@@ B.3 Gaussian memorization upper bound
where ()($\star$)( ⋆ ) is due to the fact that $h\perp\!\!\!\perp\theta\,|\,X_{1:n}$.
To construct the predictor we introduce some notation. Given a vector $v\in\mathbb{R}^{d}$, we denote by $v^{[1:\ell]}:=(v_{1},\dots,v_{\ell},0,\dots,0)\in\mathbb{R}^{d}$ its projection onto the first $\ell$ coordinates embedded in $\mathbb{R}^{d}$. Let $\hat{X}:=\frac{1}{n}\sum_{i=1}^{n}X_{i}$ be the dataset’s empirical average, and note that
$\hat{X}\sim\mathcal{N}\left(\lambda\theta,\frac{1-\lambda^{2}}{n}I_{d}\right)~{}.$ (4)
For $\hat{X}^{[1:\ell]}=(\hat{X}_{1},\dots,\hat{X}_{\ell},0,\dots,0)\in\mathbb{R}^{d}$, let $q_{k}(\hat{X}^{[1:\ell]})\in\mathbb{R}^{d}$ be a quantization of $\hat{X}^{[1:\ell]}$ that stores only $k$ bits of $\hat{X}^{[1:\ell]}$ in each coordinate’s binary expansion, allowing up to magnitude of $O(\log(d/n))$. We consider the algorithm $\mathcal{A}$ that returns the linear classifier
@@
We will argue that for suitable choices of
$\ell=\Theta(d/n)~{},~{}~{}~{}k=\Theta\left(\log(\tfrac{d}{n})\right)~{},~{}~{}~{}t=\Theta(\sqrt{d/n})~{},$
this classifier satisfies the desired properties. We note that this classifier is described by $\ell k=O(d\log(\tfrac{d}{n})/n)$ bits, so indeed Eq. (3) holds, proving the memorization upper bound.
It remains to show that $h$ is a high accuracy classifier. To bound the classifier’s error, recall Eq. (4) and note that on one hand for the null case $X\sim\mathcal{P}_{0}=\mathcal{N}(0_{d},I_{d}):$
$\langle X^{[1:\ell]},\hat{X}^{[1:\ell]}\rangle$ $=\langle X^{[1:\ell]},\lambda\theta^{[1:\ell]}\rangle+\left\langle X^{[1:\ell]},\mathcal{N}\left(0_{\ell},\frac{1-\lambda^{2}}{n}I_{\ell}\right)\right\rangle$
$=\mathcal{N}(0,\lambda^{2}\|\theta^{[1:\ell]}\|^{2}+\|g\|^{2})~{},~{}~{}~{}g\sim\mathcal{N}\left(0_{\ell},\frac{1-\lambda^{2}}{n}I_{\ell}\right)$
where equality is in the distributional sense. Hence,
$\mathop{\mathbb{E}}\langle X^{[1:\ell]},\hat{X}^{[1:\ell]}\rangle=0~{},$
and applying the general Hoeffding’s inequality [Vershynin, 2018, Theorem 2.6.2] ensures that for some absolute constants $c_{1},c_{2}>0:$
$\Pr\left[\langle X^{[1:\ell]},\hat{X}^{[1:\ell]}\rangle\geq\frac{t}{2}\right]$ $\leq\exp\left(-\frac{c_{1}t^{2}}{\lambda^{2}\|\theta^{[1:\ell]}\|^{2}+\|g\|^{2}}\right)$
$\overset{(1)}{\leq_{p}}\exp\left(-\frac{c_{2}t^{2}}{\ell/d^{1/2}+\ell/n}\right)$
$\overset{(2)}{\leq}0.001~{},$ (5)
where $\overset{(1)}{\leq_{p}}$ holds with high probability (say, at least $0.999$) for suitable $c_{2}$ since $\|\theta^{[1:\ell]}\|^{2}\lesssim\ell$ and $\|g\|^{2}\lesssim\ell\cdot\frac{1-\lambda^{2}}{n}\leq\frac{\ell}{n}$, and $\overset{(2)}{\leq}$ holds for a suitable assignment of $t,\ell$ as above, under our assumption that $d\gtrsim n^{2}$.
It remains to further bound the error induced by quantization. Note that a standard bound on the maximum of independent normal variables, each of the first $\ell$ coordinates of $\hat{X}^{[1:\ell]}$ is bounded by $\leq 1000\sqrt{2\log\ell}$ with probability at least $0.999$. Since we are storing the bits allowing up to a magnitude of $O(\log(d/n))$, under this probable event the quantization error is only due to the bits erased at the right-end of the binary expansion. It therefore holds that
$\langle X^{[1:\ell]},q_{k}(\hat{X}^{[1:\ell]})\rangle$ $=\langle X^{[1:\ell]},\hat{X}^{[1:\ell]}\rangle+\langle X^{[1:\ell]},q_{k}(\hat{X}^{[1:\ell]})-\hat{X}^{[1:\ell]}\rangle$
$\leq\langle X^{[1:\ell]},\hat{X}^{[1:\ell]}\rangle+\|X^{[1:\ell]}\|\cdot\|q_{k}(\hat{X}^{[1:\ell]})-\hat{X}^{[1:\ell]}\|$
[…]
$\leq\langle X^{[1:\ell]},\hat{X}^{[1:\ell]}\rangle+\|X^{[1:\ell]}\|\cdot\sqrt{\ell}2^{\log_{2}O(\sqrt{\log\ell})-k}$
$\leq\langle X^{[1:\ell]},\hat{X}^{[1:\ell]}\rangle+\|X^{[1:\ell]}\|\cdot\sqrt{\ell\log\ell}\cdot 2^{-k}~{},$
and further applying a Gaussian norm bound that ensures that $\Pr[\|X^{[1:\ell]}\|\geq\widetilde{C}\sqrt{\ell}]<0.0001$ for absolute constant $\widetilde{C}>0$, we see that the additional term is negligible for our setting of $k$. Union bounding with Eq. (5) and overall obtain
$\Pr_{X\sim\mathcal{P}_{0}}\left[h(X)=0\right]=\Pr_{X\sim\mathcal{P}_{0}}[\langle X,q_{k}(\hat{X}^{[1:\ell]})\rangle<t]>0.995~{}.$ (6)
As to $X\sim\mathcal{P}_{\theta}=\mathcal{N}(\lambda\theta,(1-\lambda^{2})I_{d})$, note that
$\mathop{\mathbb{E}}\langle X^{[1:\ell]},\hat{X}^{[1:\ell]}\rangle=\lambda^{2}\mathop{\mathbb{E}}\|\theta^{[1:\ell]}\|^{2}=\lambda^{2}\ell=C^{2}\cdot\Theta\left(t\right)~{}.$
Thus, similar concentration and quantization arguments as in the null case yields for sufficiently large absolute constant $C>0$ (in the problem definition):
$\Pr_{X\sim\mathcal{P}_{\theta}}\left[h(X)=1\right]=\Pr_{X\sim\mathcal{P}_{\theta}}[\langle X^{[1:\ell]},q_{k}(\hat{X}^{[1:\ell]})\rangle>t]>0.995~{}.$ (7)
By union bounding Eq. (6) and Eq. (7) we get overall that
$\mathrm{err}(\mathcal{A},\mathbf{P}_{\mathrm{G}})<0.01~{},$
which completes the proof.
@@ Appendix C Proof of Boolean Clustering Application
In this section, we prove Theorem 4.3.
@@ C.1 Boolean sample complexity $(n=1)$
To see why this predictor suffers from error at most $0.01$, first note that for $X\sim\mathcal{P}_{0}:~{}\mathop{\mathbb{E}}\langle X_{1},X\rangle=0$, and by Hoeffding’s inequality:
@@
$\mathop{\mathbb{E}}\langle X_{1},X\rangle=\sum_{i=1}^{d}\mathop{\mathbb{E}}[(X_{1})_{i}\cdot(X)_{i}]=d\cdot\mathop{\mathbb{E}}[(X_{1})_{1}]\cdot\mathop{\mathbb{E}}[(X)_{1}]~{}.$
Note that on one hand $\mathop{\mathbb{E}}[(X)_{1}]=\lambda\theta_{1}$, or equivalently $\theta_{1}\mathop{\mathbb{E}}[(X)_{1}]=\lambda$. Plugging this into the equation above, we see that
$\mathop{\mathbb{E}}\langle X_{1},X\rangle=d\theta_{1}^{2}\mathop{\mathbb{E}}[(X_{1})_{1}]\cdot\mathop{\mathbb{E}}[(X)_{1}]=d\lambda^{2}=C^{2}\sqrt{d}~{}.$
A similar calculation also shows that $\text{Var}\langle X_{1},X\rangle=d\text{Var}((X_{1})_{1})\text{Var}((X)_{1})\leq d$, following from Popoviciu’s variance bound. Therefore, Chebyshev’s inequality yields
$\Pr[\langle X_{1},X\rangle\leq(C^{2}-\sqrt{200})\sqrt{d}]\leq\frac{1}{200}~{}.$ (9)
Assuming $C$ is sufficiently large so that $C^{2}-\sqrt{200}>\sqrt{2\log(200)}$ we see by union bounding over Eq. (8) and Eq. (9) that the defined predictor $h$ indeed classifies between samples from $\mathcal{P}_{\theta}$ and $\mathcal{P}_{0}$ with error at most $\frac{1}{200}+\frac{1}{200}=\frac{1}{100}$, as claimed.
@@ C.2 Boolean memorization lower bound
Let $\lambda\in(0,1/10)$ and $n\in\mathbb{N}$ be such that $n<\frac{1}{10\lambda^{2}}$. Let $X$ be a product distribution over {±$\{\pm 1\}^{n}$ where for each coordinate $i$ it holds that $\Pr[X_{i}=1]=\frac{1+\lambda}{2}$. Similarly let $Y$ be a product distribution over {±$\{\pm 1\}^{n}$ where for each $i\in[n]$ it holds that $\Pr[Y_{i}=1]=\frac{1-\lambda}{2}$. Then, for every $\delta\in(0,1)$ and $\rho=\sqrt{8n\log(1/\delta)}\lambda$, there exists a pair of distributions $\mathbf{A},\mathbf{B}$ over {±$\{\pm 1\}^{n}$ such that
@@
Here, $p\mathbf{A}+(1-p)\mathbf{B}$ denotes the mixture of two distributions with weights $p$ and $(1-p)$.
@@
In light of Proposition C.3, to prove Lemma C.1, it suffices to show that the distributions $X,Y$ in the statement are $(\rho,\delta)$-indistinguishable.
In the language of differential privacy, $X$ and $Y$ can be understood as the $n$-wise composition of the randomized response mechanism, [^4] Recall that the $(\varepsilon,0)$-DP randomized response (RR) mechanism receives an input $b\in\{\pm 1\}$, and outputs a random bit $\hat{b}$ such that $\Pr[\hat{b}=b]=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}$. where each individual RR mechanism enjoys $(\lambda+o(1),0)$-DP. The advanced composition theorem of differential privacy then says the following:
@@
Suppose $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ are $n$ $(\varepsilon,0)$-DP algorithms operating on $b\in\{\pm 1\}$ and $n<\frac{1}{5\varepsilon^{2}}$. Then, their composition $(\mathcal{A}_{1}\circ\dots\circ\mathcal{A}_{n})$ satisfies $(\sqrt{4n\log(1/\delta)}\varepsilon,\delta)$-DP for every $\delta\in(0,1)$. In other words, $(\mathcal{A}_{1}\circ\dots\circ\mathcal{A}_{n})(-1)$ and $(\mathcal{A}_{1}\circ\dots\circ\mathcal{A}_{n})(+1)$ are $(\sqrt{4n\log(1/\delta)}\varepsilon,\delta)$-indistinguishable.
@@
We are ready to establish the second item of Theorem 4.3. Recall the setting of $\mathbf{P}_{\mathrm{B}}$: a random $\theta\sim\{\pm 1\}^{d}$ is drawn. The input distribution $\mathcal{P}_{\theta}$ is the binary symmetric channel $\mathrm{Bin}_{\frac{1-\lambda}{2}}(\theta)$. Namely, to draw $y\sim\mathcal{P}_{\theta}$, for each coordinate $i\in[d]$ set $y_{i}=\theta_{i}$ with probability $\frac{1+\lambda}{2}$, and set $y_{i}=-\theta_{i}$ otherwise. Next, we verify that the two conditions of Theorem 3.7 are met for $\mathbf{P}_{\mathrm{B}}$.
@@
Data generation SDPI. Conditioning on $\theta\in\{\pm 1\}^{d}$, for each $i\in[d]$, the $i$-th coordinates of $X_{1:n}$ are $n$ independent draws from a Bernoulli distribution of bias $\frac{1}{2}+\lambda\cdot\theta_{i}$. As such, we apply Lemma C.1 with some $\delta>0$ to be specified. Namely, we draw a single $Z_{i}$ from $\mathrm{Ber}(\frac{1}{2}+\xi\theta_{i})$ with $\xi=\sqrt{8n\log(1/\delta)}\cdot\lambda$, and post-process $Z_{i}$ as in Lemma C.1, to obtain a collection of $n$ Boolean variables $(\tilde{X}_{1,i},\dots,\tilde{X}_{n,i})$, such that (over the randomness of $Z_{i}$ and the post-processing) the TV distance between $(X_{1,i},\dots,X_{n,i})$ and $(\tilde{X}_{1,i},\dots,\tilde{X}_{n,i})$ is at most $\delta$. We run the same argument for each coordinate $i\in[d]$. Overall, we have shown how to post-process a random variable $Z\sim\mathrm{Bin}_{\frac{1-\xi}{2}}(\theta)$ to obtain a sequence of $n$ samples $\tilde{X}_{1:n}$ such that the TV distance between $\tilde{X}_{1:n}$ and $X_{1:n}$ is bounded by $\delta d$. Using Item 1 of Theorem 3.8, this implies that the pair $(\theta,X_{1:n})$ satisfies $(\xi^{2},\delta d)$-SDPI where $\xi=\sqrt{8n\log(d/\delta)}\cdot\lambda$ and $\delta\in(0,1)$ can be arbitrarily chosen.
@@
Test/train SDPI. We now verify the second condition. Consider the “$Z$” random variable we have introduced. We have shown a post-processing mapping $\Phi^{\mathrm{train}}$ such that $\Phi^{\mathrm{train}}(Z)$ is close to $X_{1:n}$ in TV distance. We can also draw $X\sim\mathcal{P}_{\theta}$ and define $\Phi^{\mathrm{test}}$ as the identity mapping. It remains to show that $Z$ and $X$ satisfy $\rho_{n}$-SDPI for a small $\rho_{n}$. Indeed, note that $\theta$ is uniform over {±$\{\pm 1\}^{d}$. Conditioned on $\theta$, we have $Z\sim\mathrm{Bin}_{\frac{1-\xi}{2}}(\theta)$ and $X\sim\mathrm{Bin}_{\frac{1-\lambda}{2}}(\theta)$, and they are independent conditioning on $\theta$. Therefore, by the symmetry of the binary symmetric channel, it follows that $Z\sim\mathrm{Bin}_{\frac{1-\lambda\xi}{2}}(X)$, which implies that $Z$ and $X$ satisfy $\rho_{n}$-SDPI with $\sqrt{\rho_{n}}=\lambda\xi=\widetilde{O}(\lambda^{2}\sqrt{n})=\widetilde{O}(\sqrt{\frac{n}{d}})$. Consequently, by Item 2 of Theorem 3.8, we conclude that the pair $(X,X_{1:n})$ satisfies $(\lambda^{2}\xi^{2},\delta d)$-SDPI, where we recall that $\xi^{2}\lambda^{2}=\widetilde{O}(\lambda^{4}n)=\widetilde{O}({\frac{n}{d}})$.
@@
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})=\Omega\left(\frac{1-\lambda^{2}n\log(d/\delta)}{\lambda^{4}n\log(d/\delta)}\left(C_{\alpha}-\delta d\log(|\mathcal{M}|/\delta)\right)\right)~{},$
which by setting $\delta=\min\{(\frac{C_{\alpha}}{d\log(|\mathcal{M}|)})^{2},\,\frac{1}{2}\}$ results in
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})=\Omega\left(\frac{1-\lambda^{2}n\log(d\log|\mathcal{M}|)}{\lambda^{4}n\log(d\log|\mathcal{M}|)}C_{\alpha}\right)~{}.$
In particular, as long as $\lambda^{2}n\leq\frac{1}{2}$, or equivalently $n\leq\sqrt{d}/C^{2}$, and if $\log|\mathcal{M}|\in\mathrm{poly}(d)$, or equivalently $|\mathcal{M}|\leq\exp(d^{\tilde{C}})$ for some absolute constant $\tilde{C}>0$, then
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})=\widetilde{\Omega}\left(\frac{1-\lambda^{2}n}{\lambda^{4}n}C_{\alpha}\right)=\widetilde{\Omega}\left(\frac{d}{n}\right)~{}.$

### v2

@@ B.3.2 Projection-based algorithms
where $(\star)$ is due to the fact that $h\perp\!\!\!\perp\theta\,|\,X_{1:n}$.
To construct the predictor we introduce some notation. Given a vector $v\in\mathbb{R}^{d}$, we denote by $v^{[1:\ell]}:=(v_{1},\dots,v_{\ell},0,\dots,0)\in\mathbb{R}^{d}$ its projection onto the first $\ell$ coordinates embedded in $\mathbb{R}^{d}$. Let $\hat{X}:=\frac{1}{n}\sum_{i=1}^{n}X_{i}$ be the dataset’s empirical average. For $\hat{X}^{[1:\ell]}=(\hat{X}_{1},\dots,\hat{X}_{\ell},0,\dots,0)\in\mathbb{R}^{d}$, let $q_{k}(\hat{X}^{[1:\ell]})\in\mathbb{R}^{d}$ be a quantization of $\hat{X}^{[1:\ell]}$ that stores only $k$ bits of $\hat{X}^{[1:\ell]}$ in each coordinate’s binary expansion, allowing up to magnitude of $O(\log(d/n))$. We consider the algorithm $\mathcal{A}$ that returns the linear classifier
@@
We claim that for some
$\ell=\frac{d}{n}~,~~~t=\Theta(\lambda^{2}\ell)=\Theta(\sqrt{d}/n)~,~~~k=\Theta\left(\log(\tfrac{d}{n})\right)~,$
this classifier satisfies the desired properties. We note that this classifier is described by $\ell k=O(d\log(\tfrac{d}{n})/n)$ bits, so Eq. (4) indeed holds, proving the memorization upper bound.
The argument that $h$ is a high accuracy classifier follows the same analysis as in Section B.1, based on the Gaussian concentration given by Lemma B.1. Indeed, noting that $\hat{X}^{[1:\ell]}\sim\mathcal{N}(\lambda\theta,\frac{1-\lambda^{2}}{n}I_{\ell})$, then simply repeating the analysis therein ensures that for $X\sim\mathcal{P}_{\theta}:~\langle X^{[1:\ell]},\hat{X}^{[1:\ell]}\rangle=\Omega(\lambda^{2}\mathop{\mathbb{E}}\|\theta^{[1:\ell]}\|)=\Omega(\lambda^{2}\ell)>t$, while in the null case, the inner product has mean zero and deviates by at most $O(\ell(\lambda^{2}+1/n))<t$ with high probability. [^4] Note that this is where the requirement that $n=O(\sqrt{d})$ is used. Furthermore, the quantization induces negligible error due to that fact that
$|\langle X^{[1:\ell]},q_{k}(\hat{X}^{[1:\ell]})\rangle-\langle X^{[1:\ell]},\hat{X}^{[1:\ell]}\rangle|\leq|\langle X^{[1:\ell]},q_{k}(\hat{X}^{[1:\ell]})-\hat{X}^{[1:\ell]}\rangle|\leq\|X^{[1:\ell]}\|\cdot\|q_{k}(\hat{X}^{[1:\ell]})-\hat{X}^{[1:\ell]}\|~,$
with the latter quantization error term going to zero as $\exp(-\Omega(k))$.
@@ Appendix C Proof of Boolean Clustering Application
In this section, we prove Theorem 4.3 assuming $C\geq(8\log(200))^{1/4}$.
@@ C.1 Boolean sample complexity $(n=1)$
To see why this predictor suffers from error at most $0.01$, first note that for $X\sim\mathcal{P}_{0}:~\mathop{\mathbb{E}}[\langle X_{1},X\rangle]=0$, and by Hoeffding’s inequality:
@@
$\mathop{\mathbb{E}}[\langle X_{1},X\rangle]=\sum_{i=1}^{d}\mathop{\mathbb{E}}[(X_{1})_{i}\cdot(X)_{i}]=d\cdot\mathop{\mathbb{E}}[(X_{1})_{1}]\cdot\mathop{\mathbb{E}}[(X)_{1}]=d\lambda^{2}=C^{2}\sqrt{d}\ ,$ (6)
and thus by Hoeffding’s inequality
$\Pr[\langle X_{1},X\rangle<\sqrt{2\log(200)d}]\leq\Pr[\langle X_{1},X\rangle<C^{2}\sqrt{d}-\sqrt{2\log(200)d}]\leq\frac{1}{200}~.$ (7)
where we used that fact that $C$ satisfies $\sqrt{2\log(200)}<C^{2}-\sqrt{2\log(200)}$. By union bounding over Eq. (5) and Eq. (7) we get that the defined predictor $h$ indeed classifies between samples from $\mathcal{P}_{\theta}$ and $\mathcal{P}_{0}$ with error at most $\frac{1}{200}+\frac{1}{200}=\frac{1}{100}$, as claimed.
@@ C.2 Boolean memorization lower bound
Let $\lambda<1$ and $n\in\mathbb{N}$ be such that $n<\frac{1}{10\lambda^{2}}$. Let $X$ be distributed according to a product distribution over $\{\pm 1\}^{n}$ where for each coordinate $i$ it holds that $\Pr[X_{i}=1]=\frac{1+\lambda}{2}$. Similarly let $Y$ be a product distribution over $\{\pm 1\}^{n}$ where for each $i\in[n]$ it holds that $\Pr[Y_{i}=1]=\frac{1-\lambda}{2}$. Then, for every $\delta\in(0,1)$ and $\rho=\sqrt{8n\log(1/\delta)}\lambda$, there exists a pair of distributions $\mathbf{A},\mathbf{B}$ over $\{\pm 1\}^{n}$ such that
@@
Here, $p\mathbf{A}+(1-p)\mathbf{B}$ denotes the mixture of two distributions with weights $p$ and (1p)(1-p).
@@
In light of Proposition C.3, to prove Lemma C.1, it suffices to show that the distributions $X,Y$ in the statement are $(\rho,\delta)$-indistinguishable, where $\rho,\delta$ are as in Lemma C.1.
In the language of differential privacy, $X$ and $Y$ can be understood as the $n$-wise parallel composition of the randomized response mechanism, [^5] Recall that the $(\varepsilon,0)$-DP randomized response (RR) mechanism receives an input $b\in\{\pm 1\}$, and outputs a random bit $\hat{b}$ such that $\Pr[\hat{b}=b]=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}$. where each individual RR mechanism enjoys $(2\lambda,0)$-DP. We can then apply the advanced composition theorem of differential privacy [Dwork et al., 2010, Kairouz et al., 2015]. We present a special case of the theorem below, which is stated in our language and suffices for our purpose.
@@
Suppose $P,Q$ are two distributions that are $(\varepsilon,0)$-indistinguishable. Let $P^{\otimes n},Q^{\otimes n}$ be the $n$-fold tensorization of $P$ and $Q$. (Namely, to sample from $P^{\otimes n}$, one independently draws $n$ samples from $P$ and list them in order). Then, $P^{\otimes n}$ and $Q^{\otimes n}$ are $(\sqrt{4n\log(1/\delta)}\varepsilon,\delta)$-indistinguishable.
@@
We are ready to establish the second item of Theorem 4.3. Recall the setting of $\mathbf{P}_{\mathrm{B}}$: a random $\theta\sim\mathcal{U}(\{\pm 1\}^{d})$ is drawn. The input distribution $\mathcal{P}_{\theta}$ is the distribution $\mathrm{BSC}_{\frac{1-\lambda}{2}}(\theta)$. Namely, to draw $y\sim\mathcal{P}_{\theta}$, for each coordinate $i\in[d]$ set $y_{i}=\theta_{i}$ with probability $\frac{1+\lambda}{2}$, and set $y_{i}=-\theta_{i}$, otherwise. Next, we verify that the two conditions of Theorem 3.7 are met for $\mathbf{P}_{\mathrm{B}}$.
@@
Data generation SDPI. Conditioning on $\theta\in\{\pm 1\}^{d}$, for each $i\in[d]$, the $i$-th coordinates of $X_{1:n}$ are $n$ independent draws from a Bernoulli distribution of bias $\frac{1}{2}+\lambda\cdot\theta_{i}$. As such, we apply Lemma C.1 with some $\delta>0$ to be specified. Namely, we define $Z_{\theta}^{\mathrm{train}}=Z$ by drawing each single $Z_{i}$ from $\mathrm{BSC}_{\frac{1-\xi}{2}}(\theta_{i})$ with $\xi=\sqrt{8n\log(1/\delta)}\cdot\lambda$, and post-process $Z_{i}$ as in Lemma C.1, to obtain a collection of $n$ Boolean variables $(\tilde{X}_{1,i},\dots,\tilde{X}_{n,i})$, such that (over the randomness of $Z_{i}$ and the post-processing) the TV distance between $(X_{1,i},\dots,X_{n,i})$ and $(\tilde{X}_{1,i},\dots,\tilde{X}_{n,i})$ is at most $\delta$. We run the same argument for each coordinate $i\in[d]$. Overall, we have shown how to post-process a random variable $Z_{\theta}^{\mathrm{train}}\sim\mathrm{BSC}_{\frac{1-\xi}{2}}(\theta)$ to obtain a sequence of $n$ samples $\tilde{X}_{1:n}$ such that the TV distance between $\tilde{X}_{1:n}$ and $X_{1:n}$ is bounded by $\delta d$. Using Item 1 of Theorem 3.8, this implies that the pair $(\theta,X_{1:n})$ satisfies $(\xi^{2},\delta d)$-SDPI where $\xi=\sqrt{8n\log(d/\delta)}\cdot\lambda$ and $\delta\in(0,1)$ can be arbitrarily chosen.
@@
Test/train SDPI. We now verify the second condition. Consider the random variable $Z$ we have introduced. We have shown a post-processing mapping $\Phi^{\mathrm{train}}$ such that $\Phi^{\mathrm{train}}(Z)$ is close to $X_{1:n}$ in TV distance. We can also draw $X\sim\mathcal{P}_{\theta}$ and define $\Phi^{\mathrm{test}}$ as the identity mapping. It remains to show that $Z$ and $X$ satisfy $\rho_{n}$-SDPI for a small $\rho_{n}$. Indeed, note that $\theta$ is uniform over $\{\pm 1\}^{d}$. Conditioned on $\theta$, we have $Z\sim\mathrm{BSC}_{\frac{1-\xi}{2}}(\theta)$ and $X\sim\mathrm{BSC}_{\frac{1-\lambda}{2}}(\theta)$, and they are independent conditioned on $\theta$. Therefore, by the symmetry of the definition of $\mathrm{BSC}$, it follows that $Z\sim\mathrm{BSC}_{\frac{1-\xi}{2}}(X)$, which implies that $Z$ and $X$ satisfy $\rho_{n}$-SDPI with $\sqrt{\rho_{n}}=\lambda\xi=O(\lambda^{2}\sqrt{n\log(d/\delta)})=O(\sqrt{n\log(d/\delta)/d})$. Consequently, by item 2 of Theorem 3.8, we conclude that the pair $(X,X_{1:n})$ satisfies $(\lambda^{2}\xi^{2},\delta d)$-SDPI, where we recall that $\xi^{2}\lambda^{2}=O(\lambda^{4}n\log(d/\delta))=O(n\log(d/\delta)/d)$.
@@
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})\geq\frac{1-8\lambda^{2}n\log(d/\delta)}{8\lambda^{4}n\log(d/\delta)}\left(C_{\alpha}-8\delta d\right)-8\delta d\log(|\mathcal{M}|/\delta d)~.$
By setting $\delta=\min\{\frac{1}{2},\,\frac{C_{\alpha}}{16d},\,(\frac{C_{\alpha}(1-8\lambda^{2}n\log(d))}{256d\lambda^{4}n\log(d)\log(|\mathcal{M}|)})^{2}\}$ so that $C_{\alpha}-8\delta d\geq C_{\alpha}/2$ and $8\delta d\log(|\mathcal{M}|/\delta d)\leq\frac{C_{\alpha}(1-8\lambda^{2}n\log(d/\delta))}{32\lambda^{4}n\log(d/\delta)}$, the bound above simplifies to
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})\geq\frac{1-8\lambda^{2}n\log(d/\delta)}{32\lambda^{4}n\log(d/\delta)}C_{\alpha}=\Omega\left(\frac{1-8\lambda^{2}n\log(d\lambda n\log|\mathcal{M}|/C_{\alpha})}{\lambda^{4}n\log(d\lambda n\log|\mathcal{M}|/C_{\alpha})}C_{\alpha}\right)~.$
[…]
$\mathrm{mem}_{n}(\mathbf{P}_{\mathrm{B}},\alpha)=\Omega\left(\frac{1-8\lambda^{2}n\log(dn/C_{\alpha})}{\lambda^{4}n\log(dn/C_{\alpha})}C_{\alpha}\right).$
For $\alpha=1/3$ we can further simplify $\mathrm{mem}_{n}(\mathbf{P}_{\mathrm{B}})=\widetilde{\Omega}\left(\frac{d}{n}\right)$.